OpenStudy (anonymous):

a hyperbola has vertices (+-5,0) and one focus (6,0) what is the standard form equation of the hyperbola

OpenStudy (phi):

Hyperbola again. Looking at http://www.clausentech.com/lchs/dclausen/algebra2/hyperbolas.htm it says that the vertices are at (h+a, k) and (h-a, k) the average of these is (h+a+ h-a, 2k)/2= (2h,2k)/2 = (h,k) (algebra again) they give us the vertices as (5,0) and (-5,0). add together: (5-5,0+0)= (0,0) divide by 2 to get the average: (0,0)/2 = (0,0) (divide by 2 didn't change anything when the center is at (0,0)

OpenStudy (phi):

so we know, so far, the center is at (0,0), the transverse axis is horizontal (it goes through the 2 vertices) the vertices are at (-5,0) and (5,0) one focus is at (6,0) the vertex is at (h+a,k). h=0 and k=0, so this is (a,0). this means a=5 (to match (5,0)